Let f (x) = 3x2 + 2 and g() = 4x – 3. Find (f o g)(-2). 245 365 -365 None of the above.
Answer #1

f(x) = 3x²+2 g(x) = 4x-3 (fog)(x) = f(g(x)) = 4(42- + 2 +2 3(4x^3) = 3 (143°+(932-233742) 3 (162²49_1242) + 2 48x²+27-72x+ 2